OpenStudy (anonymous):

A cylinder has a radius of 14 m and a height of 7 m. What is the approximate surface area of the cylinder? Use 3.14 for π. A. 1846 m2 B. 3617 m2 C. 7235 m2 D. 9847 m2

OpenStudy (rhystic):

same equation A = 2*Areaofcircle + perimeterofcircle*height. the area of the circle is pi*r^2 and perimter is 2pi*r

OpenStudy (rhystic):

you should remember the prism area equation because it works with all various shaped prisms whether it be a triangular block a cylinder etc. A = 2*base area + p * h where p is the perimeter of the base and h is the length from one base end to the other. Picture a cyclinder it has two circles and between the two circles it many circles. the two circles are your base so you find the base area of one multiply by 2 and then find the perimeter and multiply by the length from one circle to the other to get the area in between the many circles. Its the same principle with all the other prism shapes.

OpenStudy (anonymous):

So would it be b?

OpenStudy (rhystic):

no you miscalculated something. show me the numbers you are inputting.

OpenStudy (anonymous):

okay i put 3.14*14^2= 615.44 + 2*3.14*14+87.92 =...Oh wait i got 2,7142

OpenStudy (rhystic):

where did the 87.92 come from? and you need to multiply the area of the circle by 2

OpenStudy (anonymous):

I meant it =87.92

OpenStudy (rhystic):

you have the correct radius ( 14) and the area of the circl is pi*r^2 which you have correct at 3.14 * 14^2 but remember you have two of those circles so you have to multiply that result by 2. After that you have the perimter of the circl which is 2*pi*r which you have correct at 2*3.14*14 but you are forgetting to multiply that by the height of 7.

OpenStudy (rhystic):

please type the new numbers you are inputting an Ill correct you if need be.

OpenStudy (anonymous):

Okay so 3.14 * 14^2... times another 2?